# EVA: An Embodied World Model for Future Video Anticipation

### v1

@@ 5.5 Qualitative analysis
In Figure 5, we show the visualization result of our meta tasks and the rethinking. In the first example, given the initial states of the robot and human instruction, we generate the motion “Move brown chip bag near white can.” However, such task planning can not be included in 16 frames. With a self-ask rethinking, EVA extends another 16 frames until the chips bag is near the white can. The third demo is an egocentric human QA. Given the previous video and Next-Step question, EVA gives the text response “Wipe hand” together with the extending video.
In Figure 6, we show the potential of using EVA as a handbook in mixture reality. Giving the observation frame and handbook instructions in green color, EVA could keep generating long-horizon videos that guide people on how to use the kettle.
We introduce extensive visualization results, including the action of the following ability in real and simulation robots, more robot Video QA results, and the large dynamic degree of EVA in egocentric human videos in Appendix A.5 and demo page.
@@ 6 Conclusion
In this paper, we introduced the Embodied Video Anticipator (EVA), a unified framework for video understanding and generation in embodied scenarios. By decomposing complex video prediction tasks into four meta-tasks and using the Embodied Video Anticipation Benchmark (EVA-Bench) for evaluation, we demonstrated significant improvements in predicting human and robot actions. EVA’s integration of a video generation model with a visual language model, along with a multi-stage pretraining paradigm, enhances its performance in real-world applications. Our extensive experiments highlight EVA’s potential to advance large-scale pre-trained models in practical prediction tasks.
Limitation In this work, we benchmark the giving embodied video prediction task. However, such benchmarks only consider the interleaved video and language. In embodied scenes, we are expecting to introduce action, sensors, and more modalities into this dataset. Moreover, it will introduce more diverse scenarios, including autonomous driving, human-like robots, etc., in the future.
@@ A.2 Latent Diffusion Model
A.3 EVA Model Architecture
We employ a Vicuna-based VLM and a 3D U-Net architecture VDM to parameterize the EVA model. The model follows the ChatUniVi structure for VLM, and a standard 3D U-Net structure, with a spatial downsampling pass followed by an upsampling pass, utilizing skip connections from the downsampling activations. This process is interleaved with 3D convolution and attention layers. The model and training hyperparameters of EVA are summarized in Table 4 & 5.
@@ A.4 VLM Training Details
A.5 Training Dataset
To enhance comprehensive understanding, reasoning, planning, and video prediction capabilities in embodied environments, we meticulously curate a comprehensive training dataset including 500K instances, termed EVA-Instruct. This dataset encompasses four tasks, each containing videos of humans, real-world robots, and simulation robots. To enhance the diversity of prompts, we employed ChatGPT-4v (OpenAI, 2024)t to generate question-answer pairs, which were then applied to different tasks. The complete EVA instruction tuning dataset comprises 500K QA pairs collected from Open-X-Embodiment (Padalkar et al., 2023), Ego4d (Grauman et al., 2022), Ego-Exo4d (Grauman et al., 2024), and CALVIN (Mees et al., 2022). The data sources are shown as Table 6. The text in these datasets can be effectively restructured into components such as subject, verb, object, location, and destination, as illustrated in 7, making it highly suitable for our task requirements. The instructions for each meta-task are as follows.
@@ C.2 EVA Score Language Metrics
GPT-4o as a Judge: Unlike traditional similarity-based methods, GPT-4o emphasizes semantic understanding. In our implementation, we format the question, model output, and reference into a prompt, as outlined in Appendix C.4, and input it into the GPT-4o evaluator. The comparison between the generated and reference answers is based on four key criteria: object, action type, location, and attribute. During the model evaluation, we observed that some generated responses, despite being semantically close to the ground truth, received low scores. Conversely, responses omitting key information occasionally received high scores. For example, the ground truth might state, “Cut out the tomato stem with a knife on the cutting board” while Qwen2-VL-7B (Wang et al., 2024) predicts, “Chop the tomato with a knife on the cutting board”. Despite high BLEU scores, the key difference between removing the stem and chopping the tomato remains significant. Therefore, using GPT-4o as a judge to score QA text pairs and model-generated responses is essential. By leveraging GPT-4o’s advanced analytical and reasoning capabilities, we can more accurately evaluate the similarity between generated and reference texts. The specific prompt is detailed in Appendix C.4.
@@ C.3 EVA Score Video Metrics
Background Consistency: This metric evaluates the coherence between the background scene in the input image and the generated video. It utilizes DINO features and carefully designed order-statistics schemes.
@@
Aesthetic Quality: This metric refers to the perception of beauty or artistic value in a work, whether it’s traditional art, design, or even generative models. It encompasses elements like composition, color, texture, and form and how these elements evoke emotional responses from the viewer (Schuhmann et al., 2022).
Goal Completion Estimation: This metric estimates the final frame from the generative model compared to the ground truth. We evaluated the CLIP (Radford et al., 2021) feature of these two frames and normalized it. The greater the GCE, the generated image is to the target.
C.4 Model Inference Prompts

### v2

@@ 6.5 Qualitative analysis
Figures 4 showcase EVA’s capabilities in task execution, rethinking, and long-horizon guidance. In one example, EVA generates the motion ”Move brown chip bag near the white can,” and through the reflection of generation, extends frames until task completion. In an egocentric human QA scenario, EVA responds to a Next-Step question with ”Wipe hand” and generates the corresponding video. Further results, including action-following in real and simulated robots, Video QA outputs, and dynamic adaptability in egocentric videos, are detailed in Appendix B. Besides, more video demonstrations are on the anonymous page at https://sites.google.com/view/icml-eva.
@@ 7 Conclusion
In conclusion, our proposed Reflection of Generation (RoG) stretagy and the Embodied Video Anticipator (EVA) model demonstrate significant advancements in video prediction for embodied scenarios. By integrating video generation with reasoning tasks like VQA, and leveraging a chunk-wise auto-regressive strategy for longer sequences, EVA achieves high-fidelity predictions and robust generalization. The introduction of EVA-Bench further enables comprehensive and fine-grained evaluation, validating the effectiveness of RoG across diverse downstream tasks. These contributions pave the way for applying large-scale pre-trained models to real-world video prediction and embodied AI applications.
Impact Statement
Our EVA and EVA-Bench provide a unified evaluation benchmark to conduct a comprehensive assessment of the World Model’s capabilities, and advance the development of World Models in Embodied Intelligence. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ A.2 Latent Diffusion Model
A.3 Ensemble of LoRA
$W_{d}=W+\Delta W_{d}=W+A_{d}B_{d}^{T}$ (5)
$g_{d}=\text{softmax}(f(\text{Task Token}))$ (6)
$\hat{W}=W+\sum_{d}g_{d}\Delta W_{d}$ (7)
A.4 Cross-Attention Generation Adapter
A.5 EVA Model Architecture
We employ a Vicuna-based VLM and a 3D U-Net architecture VDM to parameterize the EVA model. The model follows the ChatUniVi structure for VLM, and a standard 3D U-Net structure, with a spatial downsampling pass followed by an upsampling pass, utilizing skip connections from the downsampling activations. This process is interleaved with 3D convolution and attention layers. The model and training hyperparameters of EVA are summarized in Table 6 & 7.
@@ A.6 VLM Training Details
A.7 Training Dataset
To enhance comprehensive understanding, reasoning, planning, and video prediction capabilities in embodied environments, we meticulously curate a comprehensive training dataset including 500K instances, termed EVA-Instruct. This dataset encompasses four tasks, each containing videos of humans, real-world robots, and simulation robots. To enhance the diversity of prompts, we employed ChatGPT-4v (OpenAI, 2024)t to generate question-answer pairs, which were then applied to different tasks. The complete EVA instruction tuning dataset comprises 500K QA pairs collected from Open-X-Embodiment (Padalkar et al., 2023), Ego4d (Grauman et al., 2022), Ego-Exo4d (Grauman et al., 2024), and CALVIN (Mees et al., 2022). The data sources are shown as Table 8. The text in these datasets can be effectively restructured into components such as subject, verb, object, location, and destination, as illustrated in 11, making it highly suitable for our task requirements. The instructions for each meta-task are as follows.
@@ Appendix B Appendix: Extensive experiments
B.1 Comparison in RT1
We compare the success rates of human evaluation tasks in RT1, following the framework of RoboDreamer[6]. Evaluation spans two groups—seen prompts and unseen prompts—and includes a comparison of AVDC, EVA without finish-thinking, and EVA.
Model Pick Object Move Object Near Object Place Object Upright Knock Object Over Open/close Place Object into Receptacle Summary
AVDC 11/16 9/12 1/2 4/4 2/8 2/8 29/50 (58%)
EVA (w/o finish thinking) 13/16 12/12 2/2 4/4 4/8 6/8 41/50 (82%)
EVA 13/16 12/12 2/2 4/4 4/8 7/8 42/50 (84%)
Table 9: The human evaluation results of RT-1 (Brohan et al., 2022). We perform human evaluation results to judge the task completion rate on seen prompts and seen cases of EVA and AVDC.
In seen tasks, we randomly selected 50 tasks from the validation set of RT1, including 6 tasks, across multiple scenes(Pick Object, Move Object Near Object, Place Object Upright, Knock Object Over, Open/close, Place Object into Receptacle). For the seen tasks in Tab. 9, AVDC has a 58% success rate, while EVA is 28% higher in total success rate. Moreover, EVA performed better in the Move Object task with a 100% success rate, showing good prompt-following ability. The Open/close task is especially hard since a few cases like ”open right fridge door” include the transparent glass door.
Model/Tasks Pick Object Move Object Near Object Place Object Upright Knock Object Over Open/close Place Object into Receptacle Summary
AVDC 2/4 1/4 1/1 1/2 1/1 1/3 7/15
EVA (w/o finish thinking) 2/4 1/4 0/1 2/2 0/1 0/3 5/15
EVA 3/4 4/4 1/1 2/2 1/1 1/3 12/15
Table 10: The human evaluation results of RT-1 (Brohan et al., 2022). We perform human evaluation results to judge the task completion rate on EVA and AVDC unseen prompts.
For unseen tasks, we start from the existing cases and manually change the subject or object of the prompt. For example, ”Place coke can into the bottom drawer” to ”Please close bottom drawer”. AVDC performance is better than EVA (w/o finish thinking) in Place, knock, and Open/Close tasks, since the trajectory is longer than EVA (w/o finish thinking) can generate simultaneously. However, EVA could fix this issue and significantly improve the success rate by extending the video.
@@ C.2 EVA Score Language Metrics
GPT-4o as a Judge: Unlike traditional similarity-based methods, GPT-4o emphasizes semantic understanding. In our implementation, we format the question, model output, and reference into a prompt, as outlined in Appendix C.5, and input it into the GPT-4o evaluator. The comparison between the generated and reference answers is based on four key criteria: object, action type, location, and attribute. During the model evaluation, we observed that some generated responses, despite being semantically close to the ground truth, received low scores. Conversely, responses omitting key information occasionally received high scores. For example, the ground truth might state, “Cut out the tomato stem with a knife on the cutting board” while Qwen2-VL-7B (Wang et al., 2024a) predicts, “Chop the tomato with a knife on the cutting board”. Despite high BLEU scores, the key difference between removing the stem and chopping the tomato remains significant. Therefore, using GPT-4o as a judge to score QA text pairs and model-generated responses is essential. By leveraging GPT-4o’s advanced analytical and reasoning capabilities, we can more accurately evaluate the similarity between generated and reference texts. The specific prompt is detailed in Appendix C.5.
@@ C.3 EVA Score Video Metrics
Goal Completion Estimation: In our benchmark, we manually labeled a frame as the goal condition for loop tasks like ”wash hands” lack clear ending frames. Robot tasks often have more defined endpoints, we use the last frame. We use DreamSim (Fu et al., 2023), and use an enhanced version based on DINOv2 (Caron et al., 2021), to compare the generated frames with the ground truth frames. A higher GCE indicates that the generated image is closer to the target.
C.4 Last Frame Comparison Feature Selection
To measure the similarity between the generated frames and the ground truth, we compare DreamSim(Fu et al., 2023), CLIP(Radford et al., 2021), and DINOv2(Caron et al., 2021). We visualize CLIP and DINOv2 features (Fig.12) to analyze perceptual differences in embodied scenes. Specifically:
CLIP Semantic Heatmap (Chefer et al., 2021), which highlights regions relevant to a text prompt, often misfocuses on irrelevant objects and fails to consistently capture task-relevant interactions. This likely occurs due to the absence of embodied data in CLIP’s training set. DINOv2, in contrast, excels at capturing fine-grained features, as evidenced by its attention to interactive objects such as robotic grippers. However, the visualization results of DINOv2 features heavily rely on manually adjusted PCA parameters. It often assigns high attention to background objects, and without proper PCA adjustments, the DINOv2 features of the final frames tend to exhibit extremely low variance, making it difficult to discern meaningful differences. In diverse embodied scenarios, ranging from egocentric tasks to robotic interactions, the PCA values must be frequently adjusted manually for different cases, which undermines the generalization ability of this metric in goal completion estimation tasks.
Ultimately, we select a version of DreamSim which is finetuned on DINOv2-based metric, for its improved alignment with human perception. It better aligns with human perceptual priorities by emphasizing foreground objects and semantic content while remaining sensitive to color and layout. This makes DreamSim particularly generalizable and efficient for evaluating goal state estimation in embodied tasks.
Moreover, we analyze the minimum and maximum DreamSim scores in EVA-Bench and use them to normalize DreamSim as our goal completion estimation (GCE) value, where a higher GCE score indicates closer alignment to the target.
Figure 12: Visualization of CLIP and DINOv2 Embeddings. The second row shows the attention heatmaps (relevancy) between the input text prompt and the corresponding image. The third row represents the final attention map from CLIP visual encoder. The fourth and fifth rows illustrate the PCA visualizations of DINOv2 embeddings, showing the top three principal components, with and without background masking, respectively. In the PCA visualizations, brighter regions represent higher values after the PCA transformation.
C.5 Model Inference Prompts